# Higher equations of motion at level 2 in Liouville CFT

### v1

@@ 2.4.2. Fusion estimates
This equality is holds in $\mathcal{C}^{\prime}$, but we know that the residue exists in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$ and the RHS also exists in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$. Hence the equality holds in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$. Lemma 2.5 gives the value of $\mathrm{Res}_{\alpha_{2,1}}\,J_{1}$, so we are done.
Case $\gamma>\sqrt{2}$.
In this case, $\alpha_{-2,1}=\alpha_{2,1}+2\gamma>Q$.
By symmetry, we can assume $|w_{2}|\;\leqslant\;|w_{1}|$. We split $\mathbb{D}^{2}$ into the regions
$\mathcal{D}_{0}:=\{|w_{1}|<e|w_{2}|\};\qquad\mathcal{D}_{1}:=\{|w_{1}|\;\geqslant\;e|w_{2}|\}.$
as $r\to 0$. At $\alpha=\alpha_{2,1}$, the exponent is $\frac{1}{2\gamma^{2}}(\gamma^{2}-2)^{2}>0$. By continuity in $\alpha$, this exponent is positive in a neighbourhood of $\alpha_{2,1}$. We have a covering $\mathcal{D}_{0}\subset\cup_{n\in\mathbb{N}}e^{-n}\mathbb{A}_{2}^{2}$, so that $\int_{\mathcal{D}_{0}}\Psi_{\alpha}(w_{1},w_{2})\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{|w_{2}|^{2}}$ is bounded by an absolutely convergent series.
for some $\xi>0$. This is integrable in a neighbourhood of $\alpha_{2,1}$, so we are done. ∎
@@ 2.4.3. Meromorphic continuation of the singular state
A similar formula holds for $\partial_{\bar{w}_{1}}\Psi_{\alpha}(w_{1})$.
@@
where $\alpha\mapsto\mathcal{R}_{\alpha}\in e^{-\beta c}\mathcal{D}(\mathcal{Q})$ admits a meromorphic extension which is regular in a neighbourhood $\alpha_{2,1}$. As a consequence, (2.3) holds.
@@
By integration by parts
@@
This formula is valid provided all the terms are absolutely convergent, which holds for $\alpha<\alpha_{1,2}$. We refer to [BW23, Section 3.3] for the careful justification.
Combining with the derivative formula (2.7), we get
@@
$=:\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3}.$
The integral $\mathcal{I}_{3}$ is regular for $\alpha<0$. To treat $\mathcal{I}_{2}$, we use again integration by parts and the derivative formula to get for all $F\in\mathcal{C}$:
@@
Applying this formula with $F$ replaced by $\nabla F(\partial_{w_{1}}G_{\partial}(w_{1},\cdot\,))\in\mathcal{C}$, we see that $\mathcal{I}_{2}$ is also regular for $\alpha<0$ (by [BW23, Proposition 3.2]).
Hence, the only term with a possible pole at $\alpha_{2,1}$ is $\mathcal{I}_{1}$. First, we symmetrise by observing
@@
The easiest term to treat is the first one. Applying again integration by parts and the derivative formula, we have
$\frac{\gamma}{2}(\alpha-\alpha_{1,2})\int_{\mathbb{D}^{2}}\mathbb{E}[\Psi_{\alpha}(w_{1},w_{2})F]\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{w_{2}}$ $=\gamma^{2}\int_{\mathbb{D}^{2}}\Psi_{\alpha}(w_{1},w_{2})\partial_{\bar{w}_{1}}G(w_{1},w_{2})\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{w_{2}}$
$\quad-\mu\gamma^{2}\int_{\mathbb{D}^{2}}\mathbb{E}[\Psi_{\alpha}(w_{1},w_{2},w_{3})F]\partial_{\bar{w}_{1}}G(w_{1},w_{3})\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{w_{2}}|\mathrm{d}w_{3}|^{2}$
@@
The integrals $\mathcal{I}_{6}$ and $\mathcal{I}_{7}$ are absolutely convergent and analytic for $\alpha<0$. For $\mathcal{I}_{4}$, we use the identity $\frac{1}{\bar{w}_{2}-\bar{w}_{1}}(\frac{1}{w_{1}\bar{w_{2}}w_{2}}-\frac{1}{w_{1}w_{2}\bar{w_{2}}})=\frac{1}{|w_{1}|^{2}|w_{2}|^{2}}$ and symmetrise, to see that
@@
This gives the leading term in the RHS of (2.8). It remains to show that all other terms are analytic in the neighbourhood of $\alpha_{2,1}$.
We start with the singularities with the factor $\frac{\bar{w}_{2}-\bar{w}_{1}}{w_{2}-w_{1}}$ appearing in (2.10). We only treat the singularity $\frac{\bar{w}_{2}-\bar{w}_{1}}{w_{2}-w_{1}}\frac{1}{w_{1}\bar{w}_{1}^{2}\bar{w}_{2}}$ since the other one can be treated similarly. Since $|\frac{\bar{w}_{2}-\bar{w}_{1}}{w_{2}-w_{1}}|=1$, it is sufficient to show that $\mathcal{I}_{2,(1,0),(2,1)}$ is absolutely convergent for all $\alpha$ in a neighbourhood of $\alpha_{2,1}$. We will need to treat the cases $\gamma<\sqrt{2}$ and $\gamma>\sqrt{2}$ separately. The method is similar to the proof of (2.4) given at the end of Section 2.4.2.
@@
Observe that $-2\gamma\alpha_{2,1}-\gamma^{2}+\frac{1}{2}(Q-\alpha_{2,1}-2\gamma)^{2}=\frac{1}{2}(\gamma-\frac{2}{\gamma})^{2}>0$, so that the exponent is positive in a neighbourhood of $\alpha_{2,1}$. Using the covering $\mathcal{D}_{0}\subset\cup_{n\in\mathbb{N}}e^{-n}\mathbb{A}_{2}^{2}$, the integral $\int_{\mathcal{D}_{0}}\mathbb{E}[\Psi_{\alpha}(w_{1},w_{2})F]\frac{\bar{w}_{2}-\bar{w}_{1}}{w_{2}-w_{1}}\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{\bar{w}_{2}}$ is then bounded by an absolutely convergent series, so that it converges and is analytic in a neighbourhood of $\alpha_{2,1}$.
@@
Finally, we need to analyse $\mathcal{I}_{5}$. First, we use the identity $\frac{1}{\bar{w}_{3}-\bar{w}_{1}}(\frac{1}{|w_{1}|^{2}}-\frac{1}{|w_{3}|^{2}})=\frac{1}{|w_{1}|^{2}\bar{w}_{3}}+\frac{\bar{w}_{3}-\bar{w}_{1}}{w_{3}-w_{1}}\frac{1}{\bar{w}_{1}|w_{3}|^{2}}$, and we symmetrise as before. Hence, it suffices to show that $\mathcal{I}_{3,(2,1,1),(0,0,0)}$ is absolutely convergent in a neighbourhood of $\alpha_{2,1}$. We will need to distinguish the cases $\gamma<1$, $\gamma\in[1,\sqrt{2})$, and $\gamma\in(\sqrt{2},2)$.
Case $\gamma<1$. In this case, [BW23, Proposition 3.2] says that $\mathcal{I}_{5}$ is analytic up to $\alpha<-\gamma+\frac{2}{3\gamma}$. In the range $\gamma<1$, we have $\alpha_{2,1}=-\frac{\gamma}{2}<-\gamma+\frac{2}{3\gamma}$, so in particular $\mathcal{I}_{5}$ is analytic in a neighbourhood of $\alpha_{2,1}$.
Case $\gamma\in[1,\sqrt{2})$. In this case, we have $\alpha_{2,1}+3\gamma\;\geqslant\;Q$, and $\alpha_{2,1}+2\gamma<Q$. We define the following subregions of $\mathbb{D}^{3}$
@@
$\mathcal{D}_{1}:=\{|w_{1}|\;\leqslant\;e^{-1}|w_{2}|,\,|w_{1}|\;\leqslant\;e^{-1}|w_{3}|\};\qquad\mathcal{D}_{1}^{\prime}:=\{|w_{1}|\;\geqslant\;e|w_{2}|,\,|w_{1}|\;\geqslant\;e|w_{3}|\}.$
To treat $\mathcal{D}_{0}$, we look at the case where all three insertions are in the annulus $r\mathbb{A}_{2}$ for some $r\in(0,1)$. By item 1. of Proposition 2.6, we have
@@
$-3\gamma\alpha_{2,1}-3\gamma^{2}+\frac{1}{2}(\alpha_{2,1}+3\gamma-Q)^{2}=-\frac{3\gamma^{2}}{2}+2(\frac{1}{\gamma}-\gamma)^{2}=\frac{\gamma^{2}}{2}+\frac{2}{\gamma^{2}}-4=(\frac{\gamma}{\sqrt{2}}-\frac{\sqrt{2}}{\gamma})^{2}-2>-2,$ (2.12)
@@
Let us turn to $\mathcal{D}_{1}$. In this region, we have $|w_{1}-w_{2}|^{-\gamma^{2}}\;\leqslant\;((1-e^{-1})|w_{2}|)^{-\gamma^{2}}$. By item 1. of Proposition 2.6, we have
@@
The same equality (2.12) shows that this is integrable in a neighbourhood of $\alpha_{2,1}$. In the region $\mathcal{D}_{1}^{\prime}$, we have $|w_{1}-w_{2}|^{-\gamma^{2}}\;\leqslant\;((1-e^{-1})|w_{1}|^{-\gamma^{2}}$, and Proposition 2.6 gives the bound
@@
$\quad\;\leqslant\;C\int_{\mathbb{D}}|w_{1}|^{-3\gamma\alpha-3\gamma^{2}+\frac{1}{2}(Q-3\gamma-\alpha)^{2}+2}\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}},$
so we arrive at the same conclusion. We can perform the same bounds in the regions
$\mathcal{D}_{2}:=\{|w_{2}|\;\leqslant\;e^{-1}|w_{1}|,\,|w_{2}|\;\leqslant\;e^{-1}|w_{3}|\};\qquad\mathcal{D}_{2}^{\prime}:=\{|w_{2}|\;\geqslant\;e|w_{1}|,\,|w_{2}|\;\geqslant\;e|w_{3}|\},$
@@
Case $\gamma\in(\sqrt{2},2)$. In this case, $\alpha_{2,1}+2\gamma-Q>0$. The estimate in the region $\mathcal{D}_{0}$ from the previous case is still valid, and the same conclusion holds for this region. We need to further decompose $\mathcal{D}_{1}$ into subregions. We introduce
@@
and we have seen that the exponent is positive in the neighbourhood of $\alpha_{2,1}$. This allows us to bound the integral over $\mathcal{D}_{10}$ by an absolutely convergent series, so that this integral is analytic in the neighbourhood of $\alpha_{2,1}$. In the region $\mathcal{D}_{11}$, we have $\Psi_{\alpha}(w_{1},w_{2},w_{3})=O(|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha-\gamma^{2}+\frac{1}{2}(Q-\alpha-2\gamma)^{2}}|w_{3}|^{-\gamma\alpha-\frac{3\gamma^{2}}{2}})$, and this is integrable on $(\mathbb{D}^{3},\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{|w_{2}|}\frac{|\mathrm{d}w_{3}|^{2}}{|w_{3}|})$ in the neighbourhood of $\alpha_{2,1}$. We can decompose similarly the regions $\mathcal{D}_{1}^{\prime}$, $\mathcal{D}_{2}$, $\mathcal{D}_{2}^{\prime}$ into subregions, and we get absolutely convergent integrals each time in a neighbourhood of $\alpha_{2,1}$. All these steps are similar to the previous one, and we leave the details to the reader. In the end, we have shown that $\mathcal{I}_{5}$ is analytic in a neighbourhood of $\alpha_{2,1}$, for all $\gamma\in(0,2)\setminus\{\sqrt{2}\}$.
This proves that (2.8) holds. Combining with Proposition 2.2, we obtain
@@
where $\mathcal{R}_{\alpha}$ denotes an analytic term in the neighbourhood of $\alpha_{2,1}$, which may vary from line to line. We can adapt the proof of the analyticity of $\mathcal{I}_{2,(1,0),(2,1)}(\alpha)$ to show that $\mathcal{I}_{2,(2,0),(0,2)}(\alpha)$ is analytic in the neighbourhood of $\alpha_{2,1}$. This shows that
@@
for some $\mathcal{R}_{\alpha}$ analytic at $\alpha_{2,1}$, and concludes the proof of (2.3). ∎

### v2

@@ 2.4.2. Fusion estimates
Arguing as in the proof of Proposition 2.3 shows the residue exists in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$ and the RHS also exists in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$. Hence the equality holds in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$. Lemma 2.5 gives the value of $\mathrm{Res}_{\alpha_{2,1}}\,J_{1}$, so we are done. ∎
@@ 2.4.3. Meromorphic continuation of the singular state
A similar formula holds for $\partial_{\bar{w}_{1}}\Psi_{\alpha}(\boldsymbol{w})$.
@@
where $\alpha\mapsto\mathcal{R}_{\alpha}\in e^{-\beta c}\mathcal{D}(\mathcal{Q})$ admits a meromorphic extension which is regular in a neighbourhood $\alpha_{2,1}$. Hence, the left-hand-side admits a meromorphic continuation in a neighbourhood of $\alpha_{2,1}$. In particular, (2.7) holds.
@@
The proof relies on the iteration of the same three-step strategy: integration by parts, derivative formula, and the so-called “symmetrisation trick”. As we will see below, applying this strategy to an integral with singularities (such as $\int_{\mathbb{D}}\Psi_{\alpha}(w)\frac{|\mathrm{d}w|^{2}}{|w|^{4}}$) allows us to express it as a linear combination of integrals with milder singularities (for which we have a better range of absolute convergence). The algorithm stops when we can say that each integral is either regular at $\alpha_{2,1}$ or has a pole there (that we can compute).
Let us exemplify this strategy on the integral with singularity $|w_{1}|^{-4}$. For $\alpha<\alpha_{1,2}$, we have by integration by parts
@@
This formula is valid provided all the terms are absolutely convergent, which holds for $\alpha<\alpha_{1,2}$. We refer to [BW23, Section 3.2] for the careful justification. Combining with the derivative formula (2.13), we get
@@
$=:\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3},$
where $\mathcal{I}_{j}$ is the $j^{\text{th}}$ line of the last equality. The integral $\mathcal{I}_{3}$ is regular for $\alpha<0$. To treat $\mathcal{I}_{2}$, we use again the integration by parts and the derivative formulas to get for all $F\in\mathcal{C}$:
@@
Applying this formula with $F$ replaced by $\nabla F(\partial_{w_{1}}G_{\partial}(w_{1},\cdot\,))\in\mathcal{C}$, we see that $\mathcal{I}_{2}$ is also regular for $\alpha<\alpha_{2,1}+\frac{1}{2\gamma}$ (by [BW23, Proposition 3.2]).
Hence, the only term with a possible pole at $\alpha_{2,1}$ is $\mathcal{I}_{1}$. Now, we introduce the symmetrisation trick. Note that $\partial_{w_{1}}G(w_{1},w_{2})=-\frac{1}{2(w_{1}-w_{2})}$ is antisymmetric under the permutation of $w_{1}$ and $w_{2}$, while $\Psi_{\alpha}(w_{1},w_{2})$ is permutation invariant. Hence,
$\int_{\mathbb{D}^{2}}\Psi_{\alpha}(w_{1},w_{2})\frac{1}{w_{2}-w_{1}}\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}^{2}}|\mathrm{d}w_{2}|^{2}=\frac{1}{2}\int_{\mathbb{D}^{2}}\Psi_{\alpha}(w_{1},w_{2})\frac{1}{w_{2}-w_{1}}\left(\frac{1}{w_{1}\bar{w}_{1}^{2}}-\frac{1}{w_{2}\bar{w}_{2}^{2}}\right)|\mathrm{d}w_{1}|^{2}|\mathrm{d}w_{2}|^{2}.$
Now, observe that
@@
This gives us three new types of singularities to study. We start with the one in $\frac{1}{w_{1}\bar{w}_{1}^{2}w_{2}}$. Applying again the integration by parts and the derivative formulas, we have
$\frac{\gamma}{2}(\alpha-\alpha_{1,2})\int_{\mathbb{D}^{2}}\mathbb{E}[\Psi_{\alpha}(w_{1},w_{2})F]\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{w_{2}}$ $=\gamma^{2}\int_{\mathbb{D}^{2}}\mathbb{E}[\Psi_{\alpha}(w_{1},w_{2})F]\partial_{\bar{w}_{1}}G(w_{1},w_{2})\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{w_{2}}$ (2.17)
$\quad-\mu\gamma^{2}\int_{\mathbb{D}^{3}}\mathbb{E}[\Psi_{\alpha}(w_{1},w_{2},w_{3})F]\partial_{\bar{w}_{1}}G(w_{1},w_{3})\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{w_{2}}|\mathrm{d}w_{3}|^{2}$
@@
The integrals $\mathcal{I}_{6}$ and $\mathcal{I}_{7}$ are absolutely convergent and analytic for $\alpha<\alpha_{2,1}+\frac{1}{2\gamma}$. For $\mathcal{I}_{4}$, we use the identity $\frac{1}{\bar{w}_{2}-\bar{w}_{1}}(\frac{1}{\bar{w_{1}}}-\frac{1}{\bar{w_{2}}})=\frac{1}{\bar{w}_{1}\bar{w}_{2}}$ and apply the symmetrisation trick, to see that
@@
This gives the leading term in the RHS of (2.14). It remains to show that all other terms are analytic in a neighbourhood of $\alpha_{2,1}$.
Now, we postpone the analysis of $\mathcal{I}_{5}$ to the end of the proof, and start to analyse the singularities with the factor $\frac{\bar{w}_{2}-\bar{w}_{1}}{w_{2}-w_{1}}$ appearing in (2.16). We only treat the singularity $\frac{\bar{w}_{2}-\bar{w}_{1}}{w_{2}-w_{1}}\frac{1}{w_{1}\bar{w}_{1}^{2}\bar{w}_{2}}$ since the other one can be treated similarly. Since $|\frac{\bar{w}_{2}-\bar{w}_{1}}{w_{2}-w_{1}}|=1$, it is sufficient to show that $\mathcal{I}_{2,(1,0),(2,1)}$ is absolutely convergent for all $\alpha$ in a neighbourhood of $\alpha_{2,1}$. We will need to treat the cases $\gamma<\sqrt{2}$ and $\gamma>\sqrt{2}$ separately. The method is similar to the proof of (2.8) given at the end of Section 2.4.2.
@@
Observe that $2\gamma\alpha_{2,1}-\gamma^{2}+\frac{1}{2}(Q-\alpha_{2,1}-2\gamma)^{2}=\frac{1}{2}(\gamma-\frac{2}{\gamma})^{2}>0$, so that the exponent is positive in a neighbourhood of $\alpha_{2,1}$. Using the covering $\mathcal{D}_{0}\subset\cup_{n\in\mathbb{N}}e^{-n}\mathbb{A}_{2}^{2}$, the integral $\int_{\mathcal{D}_{0}}\mathbb{E}[\Psi_{\alpha}(w_{1},w_{2})F]\frac{\bar{w}_{2}-\bar{w}_{1}}{w_{2}-w_{1}}\frac{|\mathrm{d}w_{1}|^{2}}{w_{1}\bar{w}_{1}^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{\bar{w}_{2}}$ is then bounded by an absolutely convergent series, so that it converges and is analytic in a neighbourhood of $\alpha_{2,1}$.
@@
Regularity of $\mathcal{I}_{5}$. Recall the definition of $\mathcal{I}_{5}$ in (2.17).
First, we do the symmetrisation trick by using the identity $\frac{1}{\bar{w}_{3}-\bar{w}_{1}}(\frac{1}{|w_{1}|^{2}}-\frac{1}{|w_{3}|^{2}})=\frac{1}{|w_{1}|^{2}\bar{w}_{3}}+\frac{w_{3}-w_{1}}{\bar{w}_{3}-\bar{w}_{1}}\frac{1}{w_{1}|w_{3}|^{2}}$. This shows that $\mathcal{I}_{5}$ is expressed as a linear combination where the singularities are bounded by $\frac{1}{|w_{1}|^{2}|w_{2}||w_{3}|}$ (and permutations). In other words, it suffices to show that $\mathcal{I}_{3,(2,1,1),(0,0,0)}$ is absolutely convergent in a neighbourhood of $\alpha_{2,1}$. We will need to distinguish the cases $\gamma<1$, $\gamma\in[1,\sqrt{2})$, and $\gamma\in(\sqrt{2},2)$.
Case $\gamma<1$. In this case, [BW23, Proposition 3.2] says that $\mathcal{I}_{5}$ is uniformly integrable up to $\alpha<-\gamma+\frac{2}{3\gamma}$. In the range $\gamma<1$, we have $\alpha_{2,1}=-\frac{\gamma}{2}<-\gamma+\frac{2}{3\gamma}$, so in particular $\mathcal{I}_{5}$ is analytic in a neighbourhood of $\alpha_{2,1}$.
Case $\gamma\in[1,\sqrt{2})$. In this case, we have $\alpha_{2,1}+3\gamma\;\geqslant\;Q$, and $\alpha_{2,1}+2\gamma<Q$.
By permutation symmetry, it is sufficient to treat the singularity $\frac{1}{|w_{1}|^{2}|w_{2}||w_{3}|}$ only. We define the following subregions of $\mathbb{D}^{3}$
@@
$\mathcal{D}_{1}:=\{|w_{1}|\;\leqslant\;e^{-1}|w_{2}|,\,|w_{1}|\;\leqslant\;e^{-1}|w_{3}|\}\qquad\text{and}\qquad\mathcal{D}_{1}^{\prime}:=\{|w_{1}|\;\geqslant\;e|w_{2}|,\,|w_{1}|\;\geqslant\;e|w_{3}|\}.$
To treat $\mathcal{D}_{0}$, we look at the case where all three insertions are in the annulus $r\mathbb{A}_{2}$ for some $r\in(0,1)$. By the first item of Proposition 2.6, we have
@@
$3\gamma\alpha_{2,1}-3\gamma^{2}+\frac{1}{2}(\alpha_{2,1}+3\gamma-Q)^{2}=-\frac{3\gamma^{2}}{2}+2(\frac{1}{\gamma}-\gamma)^{2}=\frac{\gamma^{2}}{2}+\frac{2}{\gamma^{2}}-4=(\frac{\gamma}{\sqrt{2}}-\frac{\sqrt{2}}{\gamma})^{2}-2>-2,$ (2.19)
@@
Let us turn to $\mathcal{D}_{1}$. In this region, we have $|w_{1}-w_{2}|^{-\gamma^{2}}\;\leqslant\;((1-e^{-1})|w_{2}|)^{-\gamma^{2}}$. By the first item of Proposition 2.6, we have
@@
Equation (2.19) shows that this is integrable in a neighbourhood of $\alpha_{2,1}$. In the region $\mathcal{D}_{1}^{\prime}$, we have $|w_{1}-w_{2}|^{-\gamma^{2}}\;\leqslant\;((1-e^{-1})|w_{1}|^{-\gamma^{2}}$, and the first item of Proposition 2.6 gives the bound
@@
$\quad\;\leqslant\;C\int_{\mathbb{D}}|w_{1}|^{-3\gamma\alpha-3\gamma^{2}+\frac{1}{2}(Q-3\gamma-\alpha)^{2}+2}\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}.$
Again, we arrive at the conclusion that the integral converges uniformly in a neighbourhood of $\alpha_{2,1}$, and is regular there. We can perform the same bounds in the regions
$\mathcal{D}_{2}:=\{|w_{2}|\;\leqslant\;e^{-1}|w_{1}|,\,|w_{2}|\;\leqslant\;e^{-1}|w_{3}|\}\qquad\text{and}\qquad\mathcal{D}_{2}^{\prime}:=\{|w_{2}|\;\geqslant\;e|w_{1}|,\,|w_{2}|\;\geqslant\;e|w_{3}|\},$
@@
Case $\gamma\in(\sqrt{2},2)$. In this case, $\alpha_{2,1}+2\gamma>Q$. The estimate in the region $\mathcal{D}_{0}$ from the previous case is still valid, and the same conclusion holds for this region. We need to further decompose $\mathcal{D}_{1}$ into subregions. We introduce
@@
and we have seen that the exponent is positive in a neighbourhood of $\alpha_{2,1}$. This allows us to bound the integral over $\mathcal{D}_{10}$ by an absolutely convergent series, so that this integral is analytic in a neighbourhood of $\alpha_{2,1}$. In the region $\mathcal{D}_{11}$, we have $\Psi_{\alpha}(w_{1},w_{2},w_{3})=O(|w_{1}|^{-\gamma\alpha}|w_{2}|^{-\gamma\alpha-\gamma^{2}+\frac{1}{2}(Q-\alpha-2\gamma)^{2}}|w_{3}|^{-\gamma\alpha-\frac{3\gamma^{2}}{2}})$, and this is integrable on $(\mathbb{D}^{3},\frac{|\mathrm{d}w_{1}|^{2}}{|w_{1}|^{2}}\frac{|\mathrm{d}w_{2}|^{2}}{|w_{2}|}\frac{|\mathrm{d}w_{3}|^{2}}{|w_{3}|})$ in a neighbourhood of $\alpha_{2,1}$. We can decompose similarly the regions $\mathcal{D}_{1}^{\prime}$, $\mathcal{D}_{2}$, $\mathcal{D}_{2}^{\prime}$ into subregions, and we get absolutely convergent integrals each time in a neighbourhood of $\alpha_{2,1}$. All these steps are similar to the previous one, and we leave the details to the reader. In the end, we have shown that $\mathcal{I}_{5}$ is analytic in a neighbourhood of $\alpha_{2,1}$, for all $\gamma\in(0,2)\setminus\{\sqrt{2}\}$.
This concludes the proof of 2.14. To end the proof of Proposition 2.7, it remains to see how this implies (2.7). Combining Proposition 2.7 with Proposition 2.2, we obtain
@@
where $\mathcal{R}_{\alpha}$ denotes an analytic term in a neighbourhood of $\alpha_{2,1}$, which may vary from line to line. It remains to show that $\mathcal{I}_{2,(2,0),(0,2)}$ is regular at $\alpha_{2,1}$; the proof is identical to that of the regularity of $\mathcal{I}_{2,(1,0),(2,1)}$ just above, and we omit it. This shows that
@@
for some $\mathcal{R}_{\alpha}$ which is regular at $\alpha_{2,1}$, and concludes the proof of (2.7). ∎